# YOLOE: Real-Time Seeing Anything

### v1

@@ Abstract
Object detection and segmentation are widely employed in computer vision applications, yet conventional models like YOLO series, while efficient and accurate, are limited by predefined categories, hindering adaptability in open scenarios. Recent open-set methods leverage text prompts, visual cues, or prompt-free paradigm to overcome this, but often compromise between performance and efficiency due to high computational demands or deployment complexity. In this work, we introduce YOLOE, which integrates detection and segmentation across diverse open prompt mechanisms within a single highly efficient model, achieving real-time seeing anything. For text prompts, we propose Re-parameterizable Region-Text Alignment (RepRTA) strategy. It refines pretrained textual embeddings via a re-parameterizable lightweight auxiliary network and enhances visual-textual alignment with zero inference and transferring overhead. For visual prompts, we present Semantic-Activated Visual Prompt Encoder (SAVPE). It employs decoupled semantic and activation branches to bring improved visual embedding and accuracy with minimal complexity. For prompt-free scenario, we introduce Lazy Region-Prompt Contrast (LRPC) strategy. It utilizes a built-in large vocabulary and specialized embedding to identify all objects, avoiding costly language model dependency. Extensive experiments show YOLOE’s exceptional zero-shot performance and transferability with high inference efficiency and low training cost. Notably, on LVIS, with 3 less training cost and 1.4 inference speedup, YOLOE-v8-S surpasses YOLO-Worldv2-S by 3.5 AP. When transferring to COCO, YOLOE-v8-L achieves 0.6 APb and 0.4 APm gains over closed-set YOLOv8-L with nearly 4 less training time. Code and models are available at https://github.com/THU-MIG/yoloe.
@@ 2 Related Work
Text-prompted detection and segmentation. Recent advancements in open-vocabulary object detection [68, 13, 76, 75, 25, 61, 74] have focused on detecting novel categories by aligning visual features with textual embeddings. Specifically, GLIP [32] unifies object detection and phrase grounding through grounded pre-training on large-scale image-text pairs, demonstrating robust zero-shot performance. DetCLIP [65] facilitates open-vocabulary learning by enriching the concepts with descriptions. Besides, Grounding DINO [37] enhances this by integrating cross-modality fusion into DINO, improving alignment between text prompts and visual representations. YOLO-World [5] further shows the potential of pretraining small detectors with open recognition capabilities based on the YOLO architecture. YOLO-UniOW [36] builds upon YOLO-World by leveraging the adaptive decision-learning strategy. Similarly, several open-vocabulary instance segmentation models [26, 11, 18, 63, 45] learn rich visual-semantic knowledge from advanced foundation models to perform segmentation on novel object categories. For example, X-Decoder [79] and OpenSeeD [71] explore both the open-vocabulary detection and segmentation tasks. APE [54] introduces a universal visual perception model that aligns and prompts all objects in image using various text prompts.
Visual-prompted detection and segmentation. While text prompts offer a generic description, certain objects can be challenging to describe with language alone, such as those requiring specialized domain knowledge. In such cases, visual prompts can guide detection and segmentation more flexibly and specifically, complementing text prompts [19, 20]. OV-DETR [67] and OWL-ViT [41] leverage CLIP encoders to process text and image prompts. MQ-Det [64] augments text queries with class-specific visual information from query images. DINOv [30] explores visual prompts as in-context examples for generic and referring vision tasks. T-Rex2 [20] integrates visual and text prompts by region-level contrastive alignment. For segmentation, based on large-scale data, SAM [23] presents a flexible and strong model that can be prompted interactively and iteratively. SEEM [80] further explores segmenting objects with more various prompt types. Semantic-SAM [31] excels in semantic comprehension and granularity detection, handling both panoptic and part segmentation tasks.

### v2

@@ Abstract
Object detection and segmentation are widely employed in computer vision applications, yet conventional models like YOLO series, while efficient and accurate, are limited by predefined categories, hindering adaptability in open scenarios. Recent open-set methods leverage text prompts, visual cues, or prompt-free paradigm to overcome this, but often compromise between performance and efficiency due to high computational demands or deployment complexity. In this work, we introduce YOLOE, which integrates detection and segmentation across diverse open prompt mechanisms within a single highly efficient model, achieving real-time seeing anything. For text prompts, we propose Re-parameterizable Region-Text Alignment (RepRTA) strategy. It refines pretrained textual embeddings via a re-parameterizable lightweight auxiliary network and enhances visual-textual alignment with zero inference and transferring overhead. For visual prompts, we present Semantic-Activated Visual Prompt Encoder (SAVPE). It employs decoupled semantic and activation branches to bring improved visual embedding and accuracy with minimal complexity. For prompt-free scenario, we introduce Lazy Region-Prompt Contrast (LRPC) strategy. It utilizes a built-in large vocabulary and specialized embedding to identify all objects, avoiding costly language model dependency. Extensive experiments show YOLOE’s exceptional zero-shot performance and transferability with high inference efficiency and low training cost. Notably, on LVIS, with 3 less training cost and 1.4 inference speedup, YOLOE-v8-S surpasses YOLO-Worldv2-S by 3.5 AP. When transferring to COCO, YOLOE-v8-L achieves 0.6 APb and 0.4 APm gains over closed-set YOLOv8-L with nearly 4 less training time. Code and models are available at here.
@@ 2 Related Work
Text-prompted detection and segmentation. Recent advancements in open-vocabulary object detection [67, 13, 75, 74, 25, 61, 73, 64] have focused on detecting novel categories by aligning visual features with textual embeddings. Specifically, GLIP [32] unifies object detection and phrase grounding through grounded pre-training on large-scale image-text pairs, demonstrating robust zero-shot performance. Grounding DINO [37] enhances this by integrating cross-modality fusion into DINO, improving alignment between text prompts and visual representations. YOLO-World [5] further shows the potential of pretraining small detectors with open recognition capabilities based on the YOLO architecture. YOLO-UniOW [36] builds upon YOLO-World by leveraging the adaptive decision-learning strategy. Similarly, several open-vocabulary instance segmentation models [26, 11, 18, 63, 45] learn rich visual-semantic knowledge from advanced foundation models to perform segmentation on novel object categories. For example, X-Decoder [78] and OpenSeeD [70] explore both the open-vocabulary detection and segmentation tasks. APE [54] introduces universal visual perception model that aligns and prompts all objects using various text prompts.
Visual-prompted detection and segmentation. While text prompts offer a generic description, certain objects can be challenging to describe with language alone, such as those requiring specialized domain knowledge. In such cases, visual prompts can guide detection and segmentation more flexibly and specifically, complementing text prompts [19, 20, 23]. OV-DETR [66] and OWL-ViT [41] leverage CLIP encoders to process text and image prompts. DINOv [30] explores visual prompts as in-context examples for generic and referring vision tasks. T-Rex2 [20] integrates visual and text prompts by region-level contrastive alignment. For segmentation, SEEM [79] explores segmenting objects with various prompt types. Semantic-SAM [31] excels in semantic comprehension and granularity detection, handling both panoptic and part segmentation.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by National Natural Science Foundation of China (Nos. 62525103, 624B2082, 62271281, 62441235).